# Physics-constrained convolutional neural networks for inverse problems in spatiotemporal partial differential equations

### v3

@@ 1 Introduction
The first inverse problem concerns error removal, which, in the literature, typically considers only the case of aleatoric noise-removal, which is a term attributed to methods that remove small, unbiased stochastic variations from the underlying solution. This has been achieved through various methods, including filtering methods (Kumar & Nachamai, 2017); proper orthogonal decomposition (Raiola et al., 2015; Mendez et al., 2017); and the use of autoencoders (Vincent et al., 2010), to name only a few. Our first demonstration of the proposed physics-constrained convolutional neural network aims, instead, to formalise and solve an inverse problem to remove large-amplitude bias (systematic error) from data; ultimately, uncovering the true solution of the partial differential equation.
@@
The second inverse problem concerns the reconstruction of solutions to partial differential equations from sparse information, which describes only a small portion of the solution and is a challenge for system identification (e.g., Brunton et al., 2016). Current methods primarily make use of convolutional neural networks due to their ability to exploit spatial correlations. The current data-driven approaches need access to the full high-resolution samples in the training, which are needed to produce a parametric mapping for the reconstruction task (e.g., Dong et al., 2014; Shi et al., 2016; Yang et al., 2019; Liu et al., 2020). In many cases, access to the high-resolution samples is limited, on which the method proposed in this paper does not rely.
@@
Common to both inverse problems that we tackle in this paper is the goal of finding a mapping from some known quantity to the corresponding solution of the governing partial differential equation. By exploiting the universal function approximation property of neural networks, it is possible to obtain numerical surrogates for function mappings or operators (Hornik et al., 1989; Zhou, 2020). For a network parameterised by weights, there exists an optimal set of weights that results in the desired mapping; the challenge being to realise these through an optimisation process. This optimisation is formulated as the minimisation of a loss function, the evaluation of which offers a distance metric to the desired solution. When considering forms of structured data, convolutional neural networks excel due to their ability to exploit spatial correlations in the data, which arise from the use of localised kernel operations.
@@
With physical systems governed by partial differential equations, the problems encountered in the absence of ground-truth high-resolution observations can be mitigated by employing a physics-constrained approach by imposing prior knowledge of the governing equations (e.g., Lagaris et al., 1998; Raissi et al., 2019; Doan et al., 2021). One such tool is that of Physics-informed neural networks (PINNs) (Raissi et al., 2019), which provide a tool for physically-motivated problems, exploiting automatic-differentiation to constrain the governing equations. Although the use of PINNs for inverse problems shows promising results for simple systems (Eivazi & Vinuesa, 2022), they remain challenging to train, and are not designed to exploit spatial correlations (Krishnapriyan et al., 2021; Grossmann et al., 2023). On the other hand, convolutional neural networks are designed to exploit spatial correlations, but they cannot naturally leverage automatic differentiation to evaluate the physical loss, as PINNs do, because they provide a mapping between states, rather than a mapping from the spatiotemporal coordinates to states as in PINNs. As such, the design of physics-informed convolutional neural networks is more challenging, and relies on finite-different approximations or differentiable solvers (Kelshaw & Magri, 2023b; Gao et al., 2021). For example, the authors of (Gao et al., 2021) show results for a steady flow field (with no temporal dynamics), which produces a mapping for stationary solutions of the Navier-Stokes equations.
@@
In this work, we propose a physics-constrained convolutional neural network with time-windowing suitable for handling a range of physically-motivated inverse problems. Imposing prior knowledge of the physics allows us to restrict the function space and use limited collocation points to find a solution. We emphasise that this is not a PINN approach Raissi et al. (2019) because we do not leverage automatic differentiation to obtain gradients of outputs with respect to inputs to constrain the physics. This network employs a time-windowing scheme to effectively compute the time-dependent residuals without resorting to recurrent-based network architectures. Realisations of the network that do not conform to the given governing equations are penalised, which ensures that the network learns to produce physical predictions.
@@ 2.1 Uncovering solutions from biased data
where ${\bm{\phi}}$ is a bias, which in practice, is spatially varying (systematic error) that can be caused by miscalibrated sensors Sciacchitano et al. (2015), or modelling assumptions Nóvoa & Magri (2022). The quantity $\bm{\zeta}(\bm{x},t)$ in Eq. (2.2) constitutes the biased dataset, which is not a solution of the PDE (2.1), i.e.
@@ 3 Convolutional neural networks
The mappings ${\bm{\eta_{\theta}}}$ and ${\bm{f_{\theta}}}$ (Eqs. (2.42.5) respectively) are design choices. For data from spatiotemporal partial differential equations, we look for a tool that: (i) can, in principle, learn any continuous functions, i.e. is a universal approximator; (ii) learn spatial correlations, as partial differential equations are defined by local operators; (iii) do not violate the conservation laws imposed by the PDE (2.1); and (iv) can handle sequential data for the temporal dynamics. In this paper, we choose convolutional neural networks (LeCun & Bengio, 1998), which are suitable tools for learning datasets of partial differential equations, from fluid mechanics Murata et al. (2020); Gao et al. (2021); Kelshaw & Magri (2023a) to heat transfer Kim & Lee (2020); Edalatifar et al. (2021). Convolutional neural networks naturally fulfil requirements (i, ii), but do not fulfil (iii, iv). In this paper, we propose the physics-constrained convolutional neural network (PC-CNN) to fulfil all requirements (i–iv). A convolutional neural network, ${\bm{k}}_{\bm{\theta}}$, is a composition of functions
@@
where ${\bm{k}}^{i}_{{\bm{\theta}}_{i}}$ denote discrete convolutional layers; $h$ is an element-wise nonlinear activation function, which increases the expressive power of the network; and $Q$ is the number of layers. The convolutional layers [^2] CNNs work on structured grids. To deal with unstructured grids and complex domains, the data can be interpolated onto a structured grid and the boundary can be padded. The extension to unstructured grids and intricate domain is beyond the scope of this paper, which is left for future work. are responsible for the discrete operation ${\bm{k}}_{\bm{\theta}}:({\bm{x}};\bm{w},\bm{b})\mapsto{\bm{w}}\ast{\bm{d}}+{\bm{b}}$, where ${\bm{\theta}}=({\bm{w}},{\bm{b}})$, $\ast$ is the convolution operation, ${\bm{x}}$ is the data, ${\bm{w}}$ are the trainable weights of the kernel, and ${\bm{b}}$ are the trainable biases. As the kernel operates locally around each pixel, information is leveraged from the surrounding grid cells. For a pedagogical explanation about CNNs the reader is referred to (Magri, 2023).
@@ 4.1.2 Losses for reconstruction from sparse information
where ${\left.\kern-1.2pt{\bm{f_{\theta}}}(\mathrel{\;\cdot\;})\vphantom{\big{|}}\right|^{{\bm{\Omega}_{L}}}}$ denotes the corestriction of ${\bm{\Omega}_{H}}$ on ${\bm{\Omega}_{L}}$. In order to find an optimal set of parameters ${\bm{\theta}}^{\ast}$, the loss is designed to regularise predictions that do not conform to the desired output. Given sparse information ${\bm{u}}({\bm{\Omega}_{L}},t)$, the data loss, $\mathcal{L}_{D}$, is defined to minimise the distance between known sparse data, ${\bm{u}}({\bm{\Omega}_{L}},t)$, and their corresponding predictions on the high-resolution grid, ${\left.\kern-1.2pt{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))\vphantom{\big{|}}\right|^{{\bm{\Omega}_{L}}}}$. Crucially, as consequence of the proposed training objective, we do not need the high-resolution field as labelled dataset, which is required with conventional super-resolution methods (Freeman et al., 2002; Dong et al., 2014; Vincent et al., 2010).
@@ 5 Synthetic data generation
We utilise a differentiable pseudospectral spatial discretisation to solve the partial differential equations of this paper Kelshaw (2022). The solution is computed on the spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{K}$, where $K$ is the number of wavenumbers. This spectral discretisation enforces periodic boundary conditions on the domain boundary, $\partial\Omega$. A solution is produced by time-integration of the dynamical system with the explicit forward-Euler scheme, in which we choose the timestep $\Delta t$ to satisfy the Courant-Friedrichs-Lewy (CFL) condition Canuto et al. (1988). The initial conditions in each case are generated using the equation
@@
As a consequence of the Nyquist-Shannon sampling criterion Canuto et al. (1988), the resolution of the spectral grid ${\bm{\hat{\Omega}}}_{k}$ places a lower bound on the spatial resolution. For a signal containing a maximal frequency $\omega_{\text{max}}$, the sampling frequency $\omega_{s}$ must satisfy the condition $\omega_{\text{max}}<\nicefrac{{\omega_{s}}}{{2}}$, therefore, we ensure that the spectral resolution satisfies $K\leq\nicefrac{{N}}{{2}}$. Violation of this condition induces spectral aliasing, in which energy content from frequencies exceeding the Nyquist limit $\nicefrac{{\omega_{s}}}{{2}}$ is fed back to the low frequencies, which amplifies energy content unphysically (Canuto et al., 1988). To prevent aliasing, in the task of uncovering solutions from biased data, we employ a spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{32\times 32}$, which corresponds to the physical grid ${\bm{\Omega}}\in\mathbb{R}^{64\times 64}$. For the task of reconstructing the solution from sparse data, we employ a high-resolution spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{35\times 35}$, which corresponds to the physical grid ${\bm{\Omega}}\in\mathbb{R}^{70\times 70}$. Approaching the Nyquist limit allows us to resolve the smallest structures possible without introducing spectral aliasing.
@@ 6 Results on uncovering solutions from biased data
We first introduce the mathematical parameterisation of the bias (i.e., systematic error). Second, we discuss numerical details and performance metrics used for the analysis. Third, we showcase results for three partial differential equations of increasing complexity: the linear convection-diffusion, nonlinear convection-diffusion (Burgers’ equation Bateman (1915); Burgers (1948)), and two-dimensional turbulent Navier-Stokes equations (the Kolmogorov flow Fylladitakis (2018)). The Navier-Stokes, which is a spatiotemporal chaotic system, will be used for the task of reconstruction from sparse information in Sec. 7. Finally, we analyse the physical consistency of physics-constrained convolutional neural network (PC-CNN) predictions for the two-dimensional turbulent flow.
@@ 6.1 Parameterisation of the bias
Errors on predictions can broadly fall into two catergories: (i) aleatoric errors, which are typically referred to as noise due to environmental and sensors’ stochastic behaviour; and (ii) epistemic uncertainties, which are typically referred to as bias. We parameterise the bias with a non-convex spatially varying function to model an epistemic error. We employ a modified Rastrigin parameterisation Rastrigin (1974), which is commonly used in non-convex optimization benchmarks
@@ 6.2 Numerical details, data, and performance metrics
We uncover the solutions of partial differential equations from biased data for three physical systems of increasing complexity. Each system is solved via the pseudospectral discretisation as described in Sec. 5, producing a solution by time-integration using the Euler-forward method with timestep of $\Delta t=5\times 10^{-3}$, which is chosen to satisfy of the Courant-Friedrichs-Lewy (CFL) condition to promote numerical stability. Model training is performed with $1024$ training samples and $256$ validation samples, which are pseudo-randomly selected from the time-domain of the solution. Each sample contains $\tau=2$ sequential time-steps. Following a hyperparameter study, the adam optimiser is employed Kingma & Ba (2015) with a learning rate of $3\times 10^{-4}$. The weighting factors for the loss, as described in Eq. 4.1, are taken as $\lambda_{P}=\lambda_{C}=10^{3}$, which are empirically determined (through a hyperparameter search) to provide stable training. Details on the network architecture can be found in A.1. Each model is trained for a total of $10^{4}$ epochs, which is chosen to provide sufficient convergence. The accuracy of prediction is quantified by the relative error on the validation dataset
@@ 6.3 The linear convection-diffusion equation
The linear convection-diffusion equation is used to describe transport phenomena Majda & Kramer (1999)
@@ 6.4 Nonlinear convection-diffusion equation
The nonlinear convection-diffusion equation, also known as Burgers’ equation Burgers (1948), is
@@ 6.4.1 Two-dimensional turbulent flow
Finally, we consider a spatiotemporally chaotic flow, which is governed by the incompressible Navier-Stokes equations evaluated on a two-dimensional domain with periodic boundary conditions, which is also known as the Kolmogorov flow Fylladitakis (2018). The partial differential equations are expressions of the mass and momentum conservation, respectively

### v4

@@ 1 Introduction
The first inverse problem concerns error removal, which, in the literature, typically considers only the case of aleatoric noise-removal, which is a term attributed to methods that remove small, unbiased stochastic variations from the underlying solution. This has been achieved through various methods, including filtering methods (Kumar and Nachamai,, 2017); proper orthogonal decomposition (Raiola et al.,, 2015; Mendez et al.,, 2017); and the use of autoencoders (Vincent et al.,, 2010), to name only a few. Our first demonstration of the proposed physics-constrained convolutional neural network aims, instead, to formalise and solve an inverse problem to remove large-amplitude bias (systematic error) from data; ultimately, uncovering the true solution of the partial differential equation.
@@
The second inverse problem concerns the reconstruction of solutions to partial differential equations from sparse information, which describes only a small portion of the solution and is a challenge for system identification (e.g., Brunton et al.,, 2016). Current methods primarily make use of convolutional neural networks due to their ability to exploit spatial correlations. The current data-driven approaches need access to the full high-resolution samples in the training, which are needed to produce a parametric mapping for the reconstruction task (e.g., Dong et al.,, 2014; Shi et al.,, 2016; Yang et al.,, 2019; Liu et al.,, 2020). In many cases, access to the high-resolution samples is limited, on which the method proposed in this paper does not rely.
@@
Common to both inverse problems that we tackle in this paper is the goal of finding a mapping from some known quantity to the corresponding solution of the governing partial differential equation. By exploiting the universal function approximation property of neural networks, it is possible to obtain numerical surrogates for function mappings or operators (Hornik et al.,, 1989; Zhou,, 2020). For a network parameterised by weights, there exists an optimal set of weights that results in the desired mapping; the challenge being to realise these through an optimisation process. This optimisation is formulated as the minimisation of a loss function, the evaluation of which offers a distance metric to the desired solution. When considering forms of structured data, convolutional neural networks excel due to their ability to exploit spatial correlations in the data, which arise from the use of localised kernel operations.
@@
With physical systems governed by partial differential equations, the problems encountered in the absence of ground-truth high-resolution observations can be mitigated by employing a physics-constrained approach by imposing prior knowledge of the governing equations (e.g., Lagaris et al.,, 1998; Raissi et al.,, 2019; Doan et al.,, 2021). One such tool is that of Physics-informed neural networks (PINNs) (Raissi et al.,, 2019), which provide a tool for physically-motivated problems, exploiting automatic-differentiation to constrain the governing equations. Although the use of PINNs for inverse problems shows promising results for simple systems (Eivazi and Vinuesa,, 2022), they remain challenging to train, and are not designed to exploit spatial correlations (Krishnapriyan et al.,, 2021; Grossmann et al.,, 2023). On the other hand, convolutional neural networks are designed to exploit spatial correlations, but they cannot naturally leverage automatic differentiation to evaluate the physical loss, as PINNs do, because they provide a mapping between states, rather than a mapping from the spatiotemporal coordinates to states as in PINNs. As such, the design of physics-informed convolutional neural networks is more challenging, and relies on finite-different approximations or differentiable solvers (Kelshaw and Magri, 2023b, ; Gao et al.,, 2021). For example, the authors of (Gao et al.,, 2021) show results for a steady flow field (with no temporal dynamics), which produces a mapping for stationary solutions of the Navier-Stokes equations.
@@
In this work, we propose a physics-constrained convolutional neural network with time-windowing suitable for handling a range of physically-motivated inverse problems. Imposing prior knowledge of the physics allows us to restrict the function space and use limited collocation points to find a solution. We emphasise that this is not a PINN approach Raissi et al., (2019) because we do not leverage automatic differentiation to obtain gradients of outputs with respect to inputs to constrain the physics. This network employs a time-windowing scheme to effectively compute the time-dependent residuals without resorting to recurrent-based network architectures. Realisations of the network that do not conform to the given governing equations are penalised, which ensures that the network learns to produce physical predictions.
@@ 2.1 Uncovering solutions from biased data
where ${\bm{\phi}}$ is a bias, which in practice, is spatially varying (systematic error) that can be caused by miscalibrated sensors Sciacchitano et al., (2015), or modelling assumptions Nóvoa and Magri, (2022). The quantity $\bm{\zeta}(\bm{x},t)$ in Eq. (2.2) constitutes the biased dataset, which is not a solution of the PDE (2.1), i.e.
@@ 3 Convolutional neural networks
The mappings ${\bm{\eta_{\theta}}}$ and ${\bm{f_{\theta}}}$ (Eqs. (2.42.5) respectively) are design choices. For data from spatiotemporal partial differential equations, we look for a tool that: (i) can, in principle, learn any continuous functions, i.e. is a universal approximator; (ii) learn spatial correlations, as partial differential equations are defined by local operators; (iii) do not violate the conservation laws imposed by the PDE (2.1); and (iv) can handle sequential data for the temporal dynamics. In this paper, we choose convolutional neural networks (LeCun and Bengio,, 1998), which are suitable tools for learning datasets of partial differential equations, from fluid mechanics Murata et al., (2020); Gao et al., (2021); Kelshaw and Magri, 2023a to heat transfer Kim and Lee, (2020); Edalatifar et al., (2021). Convolutional neural networks naturally fulfil requirements (i, ii), but do not fulfil (iii, iv). In this paper, we propose the physics-constrained convolutional neural network (PC-CNN) to fulfil all requirements (i–iv). A convolutional neural network, ${\bm{k}}_{\bm{\theta}}$, is a composition of functions
@@
where ${\bm{k}}^{i}_{{\bm{\theta}}_{i}}$ denote discrete convolutional layers; $h$ is an element-wise nonlinear activation function, which increases the expressive power of the network; and $Q$ is the number of layers. The convolutional layers [^2] CNNs work on structured grids. To deal with unstructured grids and complex domains, the data can be interpolated onto a structured grid and the boundary can be padded. The extension to unstructured grids and intricate domain is beyond the scope of this paper, which is left for future work. are responsible for the discrete operation ${\bm{k}}_{\bm{\theta}}:({\bm{x}};\bm{w},\bm{b})\mapsto{\bm{w}}\ast{\bm{d}}+{\bm{b}}$, where ${\bm{\theta}}=({\bm{w}},{\bm{b}})$, $\ast$ is the convolution operation, ${\bm{x}}$ is the data, ${\bm{w}}$ are the trainable weights of the kernel, and ${\bm{b}}$ are the trainable biases. As the kernel operates locally around each pixel, information is leveraged from the surrounding grid cells. For a pedagogical explanation about CNNs the reader is referred to (Magri,, 2023).
@@ 4.1.2 Losses for reconstruction from sparse information
where ${\left.\kern-1.2pt{\bm{f_{\theta}}}(\mathrel{\;\cdot\;})\vphantom{\big{|}}\right|^{{\bm{\Omega}_{L}}}}$ denotes the corestriction of ${\bm{\Omega}_{H}}$ on ${\bm{\Omega}_{L}}$. In order to find an optimal set of parameters ${\bm{\theta}}^{\ast}$, the loss is designed to regularise predictions that do not conform to the desired output. Given sparse information ${\bm{u}}({\bm{\Omega}_{L}},t)$, the data loss, $\mathcal{L}_{D}$, is defined to minimise the distance between known sparse data, ${\bm{u}}({\bm{\Omega}_{L}},t)$, and their corresponding predictions on the high-resolution grid, ${\left.\kern-1.2pt{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))\vphantom{\big{|}}\right|^{{\bm{\Omega}_{L}}}}$. Crucially, as consequence of the proposed training objective, we do not need the high-resolution field as labelled dataset, which is required with conventional super-resolution methods (Freeman et al.,, 2002; Dong et al.,, 2014; Vincent et al.,, 2010).
@@ 5 Synthetic data generation
We utilise a differentiable pseudospectral spatial discretisation to solve the partial differential equations of this paper Kelshaw, (2022). The solution is computed on the spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{K}$, where $K$ is the number of wavenumbers. This spectral discretisation enforces periodic boundary conditions on the domain boundary, $\partial\Omega$. A solution is produced by time-integration of the dynamical system with the explicit forward-Euler scheme, in which we choose the timestep $\Delta t$ to satisfy the Courant-Friedrichs-Lewy (CFL) condition Canuto et al., (1988). The initial conditions in each case are generated using the equation
@@
As a consequence of the Nyquist-Shannon sampling criterion Canuto et al., (1988), the resolution of the spectral grid ${\bm{\hat{\Omega}}}_{k}$ places a lower bound on the spatial resolution. For a signal containing a maximal frequency $\omega_{\text{max}}$, the sampling frequency $\omega_{s}$ must satisfy the condition $\omega_{\text{max}}<\nicefrac{{\omega_{s}}}{{2}}$, therefore, we ensure that the spectral resolution satisfies $K\leq\nicefrac{{N}}{{2}}$. Violation of this condition induces spectral aliasing, in which energy content from frequencies exceeding the Nyquist limit $\nicefrac{{\omega_{s}}}{{2}}$ is fed back to the low frequencies, which amplifies energy content unphysically (Canuto et al.,, 1988). To prevent aliasing, in the task of uncovering solutions from biased data, we employ a spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{32\times 32}$, which corresponds to the physical grid ${\bm{\Omega}}\in\mathbb{R}^{64\times 64}$. For the task of reconstructing the solution from sparse data, we employ a high-resolution spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{35\times 35}$, which corresponds to the physical grid ${\bm{\Omega}}\in\mathbb{R}^{70\times 70}$. Approaching the Nyquist limit allows us to resolve the smallest structures possible without introducing spectral aliasing.
@@ 6 Results on uncovering solutions from biased data
We first introduce the mathematical parameterisation of the bias (i.e., systematic error). Second, we discuss numerical details and performance metrics used for the analysis. Third, we showcase results for three partial differential equations of increasing complexity: the linear convection-diffusion, nonlinear convection-diffusion (Burgers’ equation Bateman, (1915); Burgers, (1948)), and two-dimensional turbulent Navier-Stokes equations (the Kolmogorov flow Fylladitakis, (2018)). The Navier-Stokes, which is a spatiotemporal chaotic system, will be used for the task of reconstruction from sparse information in Sec. 7. Finally, we analyse the physical consistency of physics-constrained convolutional neural network (PC-CNN) predictions for the two-dimensional turbulent flow.
@@ 6.1 Parameterisation of the bias
Errors on predictions can broadly fall into two catergories: (i) aleatoric errors, which are typically referred to as noise due to environmental and sensors’ stochastic behaviour; and (ii) epistemic uncertainties, which are typically referred to as bias. We parameterise the bias with a non-convex spatially varying function to model an epistemic error. We employ a modified Rastrigin parameterisation Rastrigin, (1974), which is commonly used in non-convex optimization benchmarks
@@ 6.2 Numerical details, data, and performance metrics
We uncover the solutions of partial differential equations from biased data for three physical systems of increasing complexity. Each system is solved via the pseudospectral discretisation as described in Sec. 5, producing a solution by time-integration using the Euler-forward method with timestep of $\Delta t=5\times 10^{-3}$, which is chosen to satisfy of the Courant-Friedrichs-Lewy (CFL) condition to promote numerical stability. Model training is performed with $1024$ training samples and $256$ validation samples, which are pseudo-randomly selected from the time-domain of the solution. Each sample contains $\tau=2$ sequential time-steps. Following a hyperparameter study, the adam optimiser is employed Kingma and Ba, (2015) with a learning rate of $3\times 10^{-4}$. The weighting factors for the loss, as described in Eq. 4.1, are taken as $\lambda_{P}=\lambda_{C}=10^{3}$, which are empirically determined (through a hyperparameter search) to provide stable training. Details on the network architecture can be found in A.1. Each model is trained for a total of $10^{4}$ epochs, which is chosen to provide sufficient convergence. The accuracy of prediction is quantified by the relative error on the validation dataset
@@ 6.3 The linear convection-diffusion equation
The linear convection-diffusion equation is used to describe transport phenomena Majda and Kramer, (1999)
@@ 6.4 Nonlinear convection-diffusion equation
The nonlinear convection-diffusion equation, also known as Burgers’ equation Burgers, (1948), is
@@ 6.4.1 Two-dimensional turbulent flow
Finally, we consider a spatiotemporally chaotic flow, which is governed by the incompressible Navier-Stokes equations evaluated on a two-dimensional domain with periodic boundary conditions, which is also known as the Kolmogorov flow Fylladitakis, (2018). The partial differential equations are expressions of the mass and momentum conservation, respectively